OpenStudy (anonymous):

what is 5/9 plus 1/12

OpenStudy (anonymous):

well you need to change the denominators

terenzreignz (terenzreignz):

Whatever you want it to be, I assume, Mr. Norris :) LOL Ever heard of getting the Least Common Denominator?

OpenStudy (anonymous):

so find the common demoninator

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

yep :) then you also multiply the numerator by the same number

OpenStudy (anonymous):

okieeee now what

terenzreignz (terenzreignz):

Okay... changing the denominator... \[\huge \frac{5}{9}\rightarrow \frac{\color{red}?}{36}\] So, if you change the denominator, you also have to change the numerator to keep it the same fraction... so you have to, kinda ask yourself, "what did I multiply to 9, to get 36?" Or, in quick mathematical terms, what is 36 divided by 9? The answer to both these questions is 4. So, you multiply the numerator (5) by 4, to get the new numerator... like so... \[\huge \frac{5}9 \rightarrow \frac{5\times 4}{36}= \frac{20}{36}\]

OpenStudy (anonymous):

is that the answer

terenzreignz (terenzreignz):

Nope... you have to do the same to the other addend, \[\huge \frac{1}{12}\] and then getting the answer will be much easier :) Good luck, Mr Norris, I know you don't need it :D

OpenStudy (anonymous):

so it 1/6

terenzreignz (terenzreignz):

Hardly :) because 5/9 + 1/12

terenzreignz (terenzreignz):

is

terenzreignz (terenzreignz):

Well, it isn't 1/6, anyway :)

OpenStudy (anonymous):

dangggggg ittt hahah what isss itt

terenzreignz (terenzreignz):

Well, to tell you the answer right away would be to insult your ability to arrive at the answer yourself, Mr. Norris :) Come on, you can do this... let's change the denominator of 1/12... \[\huge \frac1{12}\rightarrow \frac{\color{blue}?}{36}\] Question is... what do you multiply to 12 so that you get 36?